two Identical resistors are connected in parallel across a battery their combined resistance is 12 ohms what is the resistance of each resistor


1
Expert's answer
2021-02-03T16:17:14-0500

For parallel connection of resistors

"1 \/ R_ {sum} = 1 \/ R_1 + 1 \/ R_2 + ... + 1 \/ R_n"

where

"R_ {sum}" - equivalent resistance

"R_n" - resistance of n-th resistor


In the case of identical resistors, i.e. "R_1 = R_2 = ... = R_n" , above formula may be simplified

"R_ {sum} = R \/ n"


Answer:

Resistance of each resistor

"R = n * R_{sum} = 2 * 12 = 24 \text { (Ohm)}"